OpenStudy (anonymous):

Is 1/3 the slope of the ordered pairs (12, 36) and (18, 54)? And, what is an equation written in slope interrcept form that represents this data?

OpenStudy (nenadmatematika):

slope=(54-36)/(18-12)=18/6=3 so 1/3 is not the slope of this line....the equation of the line is: y-36=3(x-12) so y=3x

OpenStudy (anonymous):

Thank you! (: